OpenStudy (unicwaan):

Help please on two questions? A particle is moving along the x-axis so that its position at t ≥ 0 is given by s(t) = (t)In(2t). Find the acceleration of the particle when the velocity is first zero. The position of a particle on the x-axis at time t, t > 0, is s(t) = ln(t) with t measured in seconds and s(t) measured in feet. What is the average velocity of the particle for 1 ≤ t ≤ e?

OpenStudy (anonymous):

take the second derivative to get the equation for the acceleration

OpenStudy (unicwaan):

That's what I'm stuck on, not exactly how the second derivative would look with ln and e

OpenStudy (unicwaan):

@satellite73

OpenStudy (anonymous):

there is no \(e\) in it

OpenStudy (anonymous):

\[s(t)=t\ln(2t)\] the derivative requires the the product rule

OpenStudy (unicwaan):

I got ln(2t) + 1 as the first derivative and 1/t as the second derivative??

OpenStudy (unicwaan):

Also my answer choices contain e so I was a little confused

OpenStudy (anonymous):

ok first derivative is what you have that is the velocity you have to set that equal to zero and solve

OpenStudy (unicwaan):

How do I get rid of the ln?

OpenStudy (anonymous):

you mean how to you solve \[\ln(2t)+1=0\]?

OpenStudy (unicwaan):

Yes

OpenStudy (anonymous):

subtract 1 first then write in equivalent exponential form

OpenStudy (anonymous):

i can show you if that is unclear

OpenStudy (unicwaan):

I understand the ln(2x) = 1 part but after that was unclear

OpenStudy (unicwaan):

-1 sorry

OpenStudy (anonymous):

actually it should be \[\ln(2t)=-1\]

OpenStudy (anonymous):

then \[2t=e^{-1}\]

OpenStudy (anonymous):

so \[t=\frac{1}{2e}\]

OpenStudy (unicwaan):

OH! Okay that makes sense!

OpenStudy (anonymous):

now your answer should be easy since as you wrote the second derivative is \(\frac{1}{t}\) and the reciprocal of \(\frac{1}{2e}\) is \(2e\)

OpenStudy (unicwaan):

Okay I understand now, thank you! and for the second one would I only need the first derivative?

OpenStudy (anonymous):

no i don't think so

OpenStudy (anonymous):

average is just the slope of the tangent line, unless i am misreading the question

OpenStudy (unicwaan):

So I just find the slope between that interval?

OpenStudy (anonymous):

yes change in y over change is x or change in s over change in t, whatever

OpenStudy (unicwaan):

Okay, thank you so much!!!